OpenStudy (anonymous):

Which expression gives the solutions of –5 + 2x² = –6x? could someone also explain how to get this answer?

OpenStudy (anonymous):

here are the expressions

OpenStudy (anonymous):

whoops one sec

OpenStudy (anonymous):

Check out the quadratic formula. For some quadratic of the form\[ax^2 + bx + c = 0\]\[x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

1. Rearrange into standard form: 2x² + 6x – 5 = 0 then apply formula .x=−b±b2−4ac−−−−−−−√2a

OpenStudy (anonymous):

is this it?

OpenStudy (anonymous):

yeah i thnk so ..

OpenStudy (anonymous):

yea that's correct now solve for x